# Linear Complexity Computation of Code Distance and Minimum Size of Trapping Sets for LDPC Codes with Bounded Treewidth

### v1

@@ III-B Compute The Size and Number of The Smallest Trapping Set with $b=0$
$g_{t}(R,Q)=g_{t}^{\prime}(R,Q),$ (15)
for all $R\in 2^{B_{t}^{c}}$ and $Q\in 2^{B_{t}^{v}}$.
Next, we provide proof for the validity of Equation (14). For each pair $(R\subset B_{t}^{c},Q\subset B_{t}^{v})$, let $S$ be any trapping set in $G_{t}$ with parameters $(R,Q)$. $S$ remains a trapping set in $G_{t^{\prime}}$, which implies that the parameters of $S$ in $G_{t^{\prime}}$ are given by
$\begin{split}&\Big{(}\Gamma_{o}(G_{t^{\prime}}[S]),S\cap B_{t^{\prime}}^{v}\Big{)}\\ &\overset{Prop.4}{=}\Big{(}\Gamma_{o}(G_{t}[S]),S\cap B_{t}^{v}\Big{)}\\ &=(R,Q)\subset(B_{t^{\prime}}^{c},B_{t^{\prime}}^{v}).\end{split}$
This implies that $f_{t}(R,Q)\leq f_{t^{\prime}}(R,Q)$. Conversely, it is also easy to verify that any trapping set in $G_{t^{\prime}}$ with parameters $(R,Q)$ retains the same parameters in $G_{t}$, which leads to $f_{t}(R,Q)\geq f_{t^{\prime}}(R,Q)$.
Note that the reduction in the domain of the functions $f_{t}$ and $g_{t}$ in equations (14) and (15) results in the loss of certain function values. Specifically, for each pair $(R^{\prime}\subset B_{t^{\prime}}^{c},Q^{\prime}\subset B_{t^{\prime}}^{v})$ with $c\in R^{\prime}$, let $S_{t^{\prime}}$ be any trapping set in $G_{t^{\prime}}$ with parameters $(R^{\prime},Q^{\prime})$. However, since $\Gamma_{o}(G_{t}[S_{t^{\prime}}])\not\subset B_{t}^{c}$, we are unable to compute the minimum size of $S_{t^{\prime}}$ using $f_{t}$.
Nevertheless, this does not affect our ability to determine the size of the smallest trapping set with $b=0$ in $G$. This is because, according to the definition of tree decomposition, $c$ is not adjacent to any variable node in $\mathcal{L}(G)\setminus\mathcal{L}(G_{t})$. Let $S=S^{\prime}\cup S_{t^{\prime}}$ be any trapping set in $G$ such that $S^{\prime}\in\mathcal{L}(G)\setminus\mathcal{L}(G_{t})$, this leads to $c\in\Gamma_{o}(G[S])$. Consequently, the trapping set $S$ cannot correspond to $b=0$, implying that whether or not we store the minimum size of $S_{t^{\prime}}$ does not affect the final result of the algorithm.
It takes $O(2^{k})$ time for each forget check node to compute $f_{t}(R,Q)$ and $g_{t}(R,Q)$ for all $(R\subset B_{t}^{c},Q\subset B_{t}^{v})$.
Join node: Suppose $t$ is a join node, and $t_{1}$ and $t_{2}$ be its two child nodes with $B_{t}=B_{t1}=B_{t2}$. Assume we have count $f_{t_{1}}(R,Q)$, $g_{t_{1}}(R,Q)$, $f_{t_{2}}(R,Q)$, and $g_{t_{2}}(R,Q)$, for all $R\in 2^{B_{t}^{c}}$ and $Q\in 2^{B_{t}^{v}}$. Furthermore, this implies that these functions share the same domain as $f_{t}(R,Q)$ and $g_{t}(R,Q)$, which is $(2^{B_{t}^{c}},2^{B_{t}^{v}})$. We can compute $f_{t}$ and $g_{t}$ using the following equations:
$f_{t}(R,Q)=\begin{cases}F_{t}(R,Q),\text{\qquad\qquad\qquad\qquad\qquad\;\;\;\; if }Q\neq\emptyset\\ \min\{F_{t}(R,Q),f_{t_{1}}(R,Q),f_{t_{2}}(R,Q)\},\text{if }Q=\emptyset\end{cases}$ (16)
$g_{t}(R,Q)=\begin{cases}G_{t}(R,Q),&\text{if }Q\neq\emptyset;\\ \mathbb{1}_{f_{t}(R,Q)=F_{t}(R,Q)}G_{t}(R,Q)\\ +\mathbb{1}_{f_{t}(R,Q)=f_{t_{1}}(R,Q)}g_{t_{1}}(R,Q)\\ +\mathbb{1}_{f_{t}(R,Q)=f_{t_{2}}(R,Q)}g_{t_{2}}(R,Q),&\text{if }Q=\emptyset,\end{cases}$ (17)
$\begin{split}&F_{t}(R,Q)=\min_{(R_{t_{1}},R_{t_{2}})}\{f_{t_{1}}(R_{t_{1}},Q)+f_{t_{2}}(R_{t_{2}},Q)-|Q|\},\end{split}$ (18)
$\begin{split}&G_{t}(R,Q)=\sum_{(R_{t_{1}},R_{t_{2}})}g_{t_{1}}(R_{t_{1}},Q)\times g_{t_{2}}(R_{t_{2}},Q),\end{split}$ (19)
@@
for all $R\in 2^{B_{t}^{c}}$ and $Q\in 2^{B_{t}^{v}}$, where (18) takes the minimum over all pairs $(R_{t_{1}},R_{t_{2}})$ satisfying $R=R_{t_{1}}\oplus R_{t_{2}}\oplus\Gamma_{o}(G_{t}[Q\cup B_{t}^{c}])$, (19) takes the sum over all pairs $(R_{t_{1}},R_{t_{2}})$ that attain the minimum value in (18) and $\mathbb{1}_{x=y}$ is the indicator function.
First, we prove that the left-hand side of (16) is greater than or equal to the right-hand side. For each pair $(R\subset B_{t}^{c},Q\subset B_{t}^{v})$, let $S$ be any trapping set in $G_{t}$ with parameters $(R,Q)$. $S$ can be regarded as the union of three disjoint sets of variable nodes:
(1) $Q=S\cap B_{t}^{v}$;
(2) $S_{1}:=(S\setminus Q)\cap(\cup_{i\in T_{t_{1}}}B_{i})$;
(3) $S_{2}:=(S\setminus Q)\cap(\cup_{i\in T_{t_{2}}}B_{i})$.
Sets $Q$, $S_{1}$, or $S_{2}$ may be empty sets and $S$ is non-empty. Let $S_{t_{1}}=S_{1}\cup Q$ and $S_{t_{2}}=S_{2}\cup Q$. When both $S_{t_{1}}$ and $S_{t_{2}}$ are non-empty sets, let $S_{t_{1}}$ have parameters $(R_{t_{1}},Q)$ in $G_{t_{1}}$, and let $S_{t_{2}}$ have parameters $(R_{t_{2}},Q)$ in $G_{t_{2}}$. According to Lemma 1, we have $R=R_{t_{1}}\oplus R_{t_{2}}\oplus\Gamma_{o}(G_{t}[Q\cup B_{t}^{c}])$ and $|S|=|S_{t_{1}}|+|S_{t_{2}}|-|Q|$. This implies that $f_{t}(R,Q)\geq F_{t}(R,Q)$.
When $S_{t_{1}}=\emptyset$ and $S_{t_{2}}\neq\emptyset$, $S\subset(\cup_{i\in T_{t_{2}}}B_{i})$ is also a trapping set in $G_{t_{2}}$, which implies that the parameters of $S$ in $G_{t_{2}}$ are given by
$\begin{split}&\Big{(}\Gamma_{o}(G_{t_{2}}[S]),S\cap B_{t_{2}}^{v}\Big{)}\\ &=\Big{(}\Gamma_{o}(G_{t_{2}}[S_{2}]),S\cap B_{t}^{v}\Big{)}\\ &\overset{Prop.3}{=}\Big{(}\Gamma_{o}(G_{t}[S_{2}\cup(\cup_{i\in T_{t_{2}}}B_{i}^{c})]),S\cap B_{t}^{v}\Big{)}\\ &\overset{Prop.6}{=}\Big{(}\Gamma_{o}(G_{t}[S_{2}]),Q\Big{)}\\ &=\Big{(}\Gamma_{o}(G_{t}[S]),Q\Big{)}\\ &=(R,Q).\end{split}$ (20)
Therefore, $f_{t}{(R,Q)}\geq f_{t_{2}}(R,Q)$.
When $S_{t_{1}}\neq\emptyset$ and $S_{t_{2}}=\emptyset$, a similar proof shows that $f_{t}(R,Q)\geq f_{t_{1}}(R,Q)$.
By combining these three cases, we can prove that the left-hand side of Equation (16) is greater than or equal to the right-hand side.
On the other hand, we prove that the left-hand side of Equation (16) is less than or equal to the right-hand side. For each pair $(R\subset B_{t}^{c},Q\subset B_{t}^{v})$, let $S_{t_{1}}$ and $S_{t_{2}}$ be any trapping sets in $G_{t_{1}}$ and $G_{t_{2}}$ with parameters $(R_{t_{1}},Q)$ and $(R_{t_{2}},Q)$, respectively, such that $R=R_{t_{1}}\oplus R_{t_{2}}\oplus\Gamma_{o}(G_{t}[Q\cup B_{t}^{c}])$. Let $S_{t}=S_{t_{1}}\cup S_{t_{2}}$. Then, according to Lemma 1, $S_{t}$ is a trapping set in $G_{t}$ with parameters $(R,Q)$. Thus, $F_{t}(R,Q)\geq f_{t}(R,Q).$ Furthermore, when $Q=\emptyset$, let $S_{t_{j}}^{\prime}$ be any trapping set with parameters $(R,\emptyset)$ in $G_{t_{j}}$, where $j=1$ and 2. Then, $S_{t_{j}}^{\prime}$ is also a trapping set in $G_{t}$. According to (20), we have that the parameters of $S_{t_{j}}^{\prime}$ in $G_{t}$ are given by $(R,Q)$. Therefore, we have $f_{t_{j}}(R,\emptyset)\geq f_{t}(R,\emptyset)$.
@@
For each pair $(R\subset B_{t}^{c},Q\subset B_{t}^{v})$, it takes $O(2^{k})$ time to obtain the minimum value in (18). Therefore, the computational complexity for each join node is $O(2^{2k})$, while for each other type of node, it is $O(2^{k})$. Given that a nice tree decomposition $(T,(B_{t})_{t\in T})$ of the Tanner graph $G$ contains $O(n)$ nodes in $T$, the overall time complexity of the algorithm is $O(n2^{2k})$. ∎
Note that during the proof process, the computation of $f_{t}$ and $g_{t}$ depends only on the functions $f_{t^{\prime}}$ and $g_{t^{\prime}}$, where $t^{\prime}$ is the child of $t$ in $T$. Therefore, when traversing the bags from the leaves to the root, the functions $f_{i}$ and $g_{i}$, where $i$ is a descendant of $t^{\prime}$, do not need to be stored.
By executing the algorithm in Theorem 2 several times, we can find an $(a,0)$-trapping set with minimum $a$ in $G$. This leads to the following corollary:
@@
Let $G$ be a Tanner graph with treewidth $tw(G)=k$, corresponding to an LDPC code of length $n$. We can find an $(a,0)$-trapping sets with minimum $a$ in $G$ in time $O(n2^{2k})$.
@@
In fact, we can first execute the algorithm in Theorem 2 to determine the minimum size $a_{0}$ of an $(a,0)$-trapping set. Then, by running the algorithm $a_{0}$ times, we can identify all $(a_{0},0)$-trapping sets. ∎
Sometimes, computing the treewidth of a given LDPC code is challenging. However, if the pathwidth of the LDPC code is known, we can derive the following corollary by applying the algorithm in Theorem 2.
@@
Let $G$ be a Tanner graph with path-width $pw(G)=k^{\prime}$, corresponding to an LDPC code of length $n$. Then, the size and the number of the smallest trapping sets with $b=0$ in $G$ can be determined in time $O(n2^{k^{\prime}})$.
@@
Let $(T,(B_{t})_{t\in T})$ be a nice tree decomposition of $G$ with width $k^{\prime}$, where $T$ is a path. In this case, $T$ contains no join nodes and consists only of leaf nodes, introduced variable nodes, forget variable nodes, introduced check nodes, and forget check nodes. According to Theorem 2, each node requires $O(2^{k^{\prime}})$ time, leading to an overall computational complexity of $O(n2^{k^{\prime}})$ for the algorithm. ∎

### v2

@@ III-B Compute The Size and Number of The Smallest Trapping Set with $b=0$
$g_{t_{i}}(L,I)=\sum_{\begin{subarray}{c}P\in\{(L,I),(L\cup\{v\},I)\}\\ f_{t_{i}}(L,I)=f_{t_{j}}(P)\end{subarray}}g_{t_{j}}(P).$ (10)
Let $S$ be a trapping set in $G$. For each pair $(L\subset\mathcal{L}^{B}_{i},I\subset\mathcal{R}^{B}_{i})$, we claim that $S$ is a $(L,I,t_{i})$-type trapping set if and only if $S$ is either a $(L,I,t_{j})$-type trapping set or a $(L\cup\{v\},I,t_{j})$-type trapping set, which leads to (9) and (10). According to $B_{t_{i}}=B_{t_{j}}\setminus\{v\}$, we have:
$\begin{split}S\subset\mathcal{L}_{j}\Leftrightarrow S\subset\mathcal{L}_{i}\end{split}$ (11)
$\begin{split}S\cap B_{t_{j}}&=S\cap(B_{t_{i}}\cup\{v\})\\ &=(S\cap B_{t_{i}})\cup(S\cap\{v\}),\end{split}$ (12)
and
$\begin{split}\Gamma^{o}_{G_{j}}(S)=\Gamma^{o}_{G_{i}}(S).\end{split}$ (13)
Clearly, based on equations (11) $\sim$ (13), the sufficiency of the claim is established. Conversely, to establish necessity, suppose $S$ is a $(L,I,t_{j})$-type trapping set. Since $v\notin L,v\in B_{t_{j}}$, and $S\cap B_{t_{j}}=L$, we have $S\cap\{v\}=\emptyset$, which implies that $S\cap B_{t_{i}}=L$. Therefore, $S$ is a $(L,I,t_{i})$-type trapping set. Similarly, suppose $S$ is a $(L\cup\{v\},I,t_{j})$-type trapping set. According to (12), $S\cap B_{t_{i}}=L$ or $S\cap B_{t_{i}}=L\cup\{v\}$. Since $v\notin B_{t_{i}}$, it follows that $S\cap B_{t_{i}}=L$, and thus $S$ is a $(L,I,t_{i})$-type trapping set.
It takes $O(2^{k})$ time for each forget variable node to compute $f_{t_{i}}(L,I)$ and $g_{t_{i}}(L,I)$ for all $(L,I)$.
introduced check node: Suppose $t_{i}$ is an introduced node with a unique child $t_{j}$, and that the corresponding bags satisfy $B_{t_{i}}=B_{t_{j}}\cup\{c\}$, where $c$ is a check node. We can compute $f_{t_{i}}$ and $g_{t_{i}}$ using the following equations for all $L\subset{\mathcal{L}^{B}_{i}}$, $I\subset{\mathcal{R}^{B}_{i}}$:
$f_{t_{i}}(L,I)=\begin{cases}f_{t_{j}}(L,I\setminus\{c\}),&\text{if }c\in I\text{, $E(c,L)$ is odd;}\\ f_{t_{j}}(L,I),&\text{if }c\notin I\text{, $E(c,L)$ is even;}\\ +\infty,&\text{otherwise},\end{cases}$ (14)
$g_{t_{i}}(L,I)=\begin{cases}g_{t_{j}}(L,I\setminus\{c\}),&\text{if }c\in I\text{, $E(c,L)$ is odd;}\\ g_{t_{j}}(L,I),&\text{if }c\notin I\text{, $E(c,L)$ is even;}\\ 0,&\text{otherwise},\end{cases}$ (15)
where $E(c,L)$ denotes the number of edges between $c$ and $L$ in $G_{i}$.
Next, we prove the correctness of (14) and (15). Let $S$ be a trapping set in $G$. For each pair $(L\subset\mathcal{L}^{B}_{i},I\subset\mathcal{R}^{B}_{i})$, we have:
$\begin{split}S\subset\mathcal{L}_{j}\Leftrightarrow S\subset\mathcal{L}_{i}\end{split}$ (16)
$\begin{split}S\cap B_{t_{j}}=S\cap B_{t_{i}}\end{split}$ (17)
and
$\begin{split}\Gamma_{G_{j}}^{o}(S)&=\Gamma_{G_{i}\setminus\{c\}}^{o}(S)\\ &\overset{prop.2}{=}\Gamma_{G_{i}}^{o}(S)\oplus\Gamma_{G[c\cup S]}^{o}(S)\\ &\overset{(a)}{=}\Gamma_{G_{i}}^{o}(S)\oplus\Gamma_{G[c\cup S]}^{o}(S\cap B_{t_{i}})\end{split}$ (18)
The validity of $(a)$ follows from the definition of treewidth: since $c$ appears for the first time in $B_{t_{i}}$, it can be adjacent only to variable nodes within $B_{t_{i}}$, implying that $E(c,S)=E(c,S\cap B_{t_{i}})$. Therefore, the following equation holds:
$\begin{split}&\Gamma_{G[c\cup S]}^{o}(S)\\ &=\Gamma_{G[c\cup S]}^{o}(S\cap B_{t_{i}})=\begin{cases}\{c\},&\text{if }E(c,S\cap B_{t_{i}})\text{ is odd}\\ \emptyset,&\text{if }E(c,S\cap B_{t_{i}})\text{ is even}\end{cases}\end{split}$ (19)
  • If $E(c,L)$ is odd and $c\in I$, according to (16)–(19), we have that $S$ is a $(L,I,t_{i})$-type trapping set if and only if $S$ is a $(L,I\setminus\{c\},t_{j})$-type trapping set. Consequently, $f_{t_{i}}(L,I)=f_{t_{j}}(L,I\setminus\{c\})$ and $g_{t_{i}}(L,I)=g_{t_{j}}(L,I\setminus\{c\})$.
  • If $E(c,L)$ is even and $c\notin I$, according to (16)–(19), we have that $S$ is a $(L,I,t_{i})$-type trapping set if and only if $S$ is a $(L,I,t_{j})$-type trapping set. Consequently, $f_{t_{i}}(L,I)=f_{t_{j}}(L,I)$ and $g_{t_{i}}(L,I)=g_{t_{j}}(L,I)$.
  • If $E(c,L)$ is odd and $c\notin I$, we claim that there does not exist an $(L,I,t_{i})$-type trapping set $S$ in $G_{i}$. Otherwise, since $E(c,L=S\cap B_{t_{i}})$ is odd, we have $E(c,S)$ is odd and $c\in\Gamma^{o}_{G_{i}}(S)$, which contradicts the assumption that $c\notin I$.
  • If $E(c,L)$ is even and $c\in I$, we claim that there does not exist an $(L,I,t_{i})$-type trapping set $S$ in $G_{i}$. Otherwise, Since $E(c,L)$ is even, we have $E(c,S)$ is even and $c\notin\Gamma^{o}_{G_{i}}(S)$, which contradicts the assumption that $c\in I$.
It takes $O(2^{k})$ time for each introduced check node to compute $f_{t_{i}}(L,I)$ and $g_{t_{i}}(L,I)$ for all $(L,I)$.
Forget check node: Suppose $t_{i}$ is a forget node with a unique child $t_{j}$, and that the corresponding bags satisfy $B_{t_{i}}=B_{t_{j}}\setminus\{c\}$, where $c$ is a check node. We can compute $f_{t_{i}}$ and $g_{t_{i}}$ using the following equations for all $L\subset{\mathcal{L}^{B}_{i}}$, $I\subset{\mathcal{R}^{B}_{i}}$:
$f_{t_{i}}(L,I)=f_{t_{j}}(L,I),$ (20)
and
$g_{t_{i}}(L,I)=g_{t_{j}}(L,I).$ (21)
Let $S$ be a trapping set in $G$. For each pair $(L\subset\mathcal{L}^{B}_{i},I\subset\mathcal{R}^{B}_{i})$, we claim that $S$ is a $(L,I,t_{i})$-type trapping set if and only if $S$ is a $(L,I,t_{j})$-type trapping set, which leads to $f_{t_{i}}(L,I)=f_{t_{j}}(L,I)$ and $g_{t_{i}}(L,I)=g_{t_{j}}(L,I)$. This claim follows from the fact that:
$\begin{split}S\subset\mathcal{L}_{j}\Leftrightarrow S\subset\mathcal{L}_{i}\end{split}$ (22)
$\begin{split}S\cap B_{t_{j}}=S\cap B_{t_{i}}\end{split}$ (23)
and
$\begin{split}\Gamma_{G_{j}}^{o}(S)&=\Gamma_{G_{i}}^{o}(S)\end{split}$ (24)
It takes $O(2^{k})$ time for each forget check node to compute $f_{t_{i}}(L,I)$ and $g_{t_{i}}(L,I)$ for all $(L,I)$.
Join node: Suppose $t_{i}$ is a join node, and $t_{j_{1}}$ and $t_{j_{2}}$ be its two child nodes with $B_{t_{i}}=B_{t_{j_{1}}}=B_{t_{j_{2}}}$. We define the following two functions:
$F_{t_{i}}(L,I)=\min_{\begin{subarray}{c}I^{\prime},I^{\prime\prime}\end{subarray}}\{f_{t_{j_{1}}}(L,I^{\prime})+f_{t_{j_{2}}}(L,I^{\prime\prime})-|L|\}.$ (25)
$\begin{split}&G_{t_{i}}(L,I)=\sum_{\begin{subarray}{c}I^{\prime},I^{\prime\prime}\end{subarray}}g_{t_{j_{1}}}(L,I^{\prime})\times g_{t_{j_{2}}}(L,I^{\prime\prime}).\end{split}$ (26)
@@
where (25) takes the minimum over all $I^{\prime},I^{\prime\prime}$ satisfying $I=I^{\prime}\oplus I^{\prime\prime}\oplus\Gamma^{o}_{G[B_{t}]}(L)$, (26) takes the sum over all $I^{\prime},I^{\prime\prime}$ that attain the minimum value in (25). We can compute $f_{t_{i}}$ and $g_{t_{i}}$ using the following equations for all $L\subset{\mathcal{L}^{B}_{i}}$, $I\subset{\mathcal{R}^{B}_{i}}$:
$f_{t_{i}}(L,I)=\begin{cases}F_{t_{i}}(L,I),\text{\qquad\qquad\qquad\qquad\qquad\;\;\;\; if }L\neq\emptyset\\ \min\{F_{t_{i}}(L,I),f_{t_{j_{1}}}(L,I),f_{t_{j_{2}}}(L,I)\},\text{if }L=\emptyset\end{cases}$ (27)
$g_{t_{i}}(L,I)=\begin{cases}G_{t_{i}}(L,I),&\text{if }L\neq\emptyset;\\ \mathbb{1}_{f_{t_{i}}(L,I)=F_{t_{i}}(L,I)}G_{t_{i}}(L,I)\\ +\sum_{t\in\{t_{j_{1}},t_{j_{2}}\}}\mathbb{1}_{f_{t_{i}}(L,I)=f_{t}(L,I)}g_{t}(L,I),&\text{if }L=\emptyset,\end{cases}$ (28)
where $\mathbb{1}_{x=y}$ is the indicator function.
Let $S$ be a trapping set in $G_{i}$, and define
$S_{0}=S\cap B_{t_{i}}$
$S_{1}=S\cap G_{t_{j_{1}}}$
$S_{2}=S\cap G_{t_{j_{2}}}$
We claim that $S$ is a $(L,I,t_{i})$-type trapping set if and only if $S_{1}$ is a $(L,I^{\prime},t_{j_{1}})$-type trapping set and $S_{2}$ is a $(L,I^{\prime\prime},t_{j_{2}})$-type trapping set where $I=I^{\prime}\oplus I^{\prime\prime}\oplus\Gamma^{o}_{G[B_{t_{i}}]}(L)$. This claim follows from the fact that:
$\begin{split}S\subset\mathcal{L}_{i},S_{1}\subset\mathcal{L}_{j_{1}},S_{2}\subset\mathcal{L}_{j_{2}}\end{split}$ (29)
$\begin{split}&S_{1}\cap B_{t_{j_{1}}}=S\cap G_{t_{j_{1}}}\cap B_{t_{j_{1}}}\\ &=S\cap B_{t_{i}}=S_{2}\cap B_{t_{j_{2}}}\end{split}$ (30)
and
$\begin{split}&\Gamma_{G_{i}}^{o}(S)\\ &=\Gamma_{G_{i}}^{o}(S_{0}\cup(S_{1}\setminus S_{0})\cup(S_{2}\setminus S_{0}))\\ &\overset{(b)}{=}\Gamma_{G_{i}}^{o}(S_{0})\oplus\Gamma_{G_{i}}^{o}(S_{1}\setminus S_{0})\oplus\Gamma_{G_{i}}^{o}(S_{2}\setminus S_{0})\\ &\overset{(c)}{=}\Gamma_{G_{i}}^{o}(S_{0})\oplus\Gamma_{G_{j_{1}}}^{o}(S_{1}\setminus S_{0})\oplus\Gamma_{G_{j_{2}}}^{o}(S_{2}\setminus S_{0})\\ &\overset{(d)}{=}\Gamma_{G_{i}}^{o}(S_{0})\oplus\Gamma_{G_{j_{1}}}^{o}(S_{1})\oplus\Gamma_{G_{j_{1}}}^{o}(S_{0})\oplus\Gamma_{G_{j_{2}}}^{o}(S_{2})\oplus\Gamma_{G_{j_{2}}}^{o}(S_{0})\end{split}$ (31)
The equality in (b) follows from the definitions of $S_{0}$, $S_{1}$, and $S_{2}$. In particular, the sets $S_{0}$, $S_{1}\setminus S_{0}$, and $S_{2}\setminus S_{0}$ are pairwise disjoint, and the result follows directly from Proposition 1. The equality in (c) is a consequence of the definition of a tree decomposition. Specifically, the set $S_{1}\setminus S_{0}=S\cap(G_{t_{j_{1}}}\setminus B_{t_{i}})$ is adjacent only to check nodes in $G_{t_{j_{1}}}$, which implies that
$\Gamma_{G_{i}}^{o}(S_{1}\setminus S_{0})=\Gamma_{G_{j_{1}}}^{o}(S_{1}\setminus S_{0})$
The equality in $d$ follows from the inclusio $S_{0}\subseteq S_{1}$, and again from Proposition 1.
According to the Proposition 2, we have
$\begin{split}\Gamma_{G_{i}}^{o}(S_{0})=\Gamma_{G_{i}\setminus G_{j_{1}}}^{o}(S_{0})\oplus\Gamma_{G_{i}\setminus G_{j_{2}}}^{o}(S_{0})\oplus\Gamma_{G[B_{t_{i}}]}^{o}(S_{0})\end{split}$ (32)
$\begin{split}\Gamma_{G_{j_{1}}}^{o}(S_{0})&=\Gamma_{G_{j_{1}}\setminus G[B_{t_{i}}]}^{o}(S_{0})\oplus\Gamma_{G[B_{t_{i}}]}^{o}(S_{0})\\ &=\Gamma_{G_{i}\setminus G_{j_{2}}}^{o}(S_{0})\oplus\Gamma_{G[B_{t_{i}}]}^{o}(S_{0})\end{split}$ (33)
$\begin{split}\Gamma_{G_{j_{2}}}^{o}(S_{0})=\Gamma_{G_{i}\setminus G_{j_{1}}}^{o}(S_{0})\oplus\Gamma_{G[B_{t_{i}}]}^{o}(S_{0})\end{split}$ (34)
Substituting these equations into (31), we obtain
$\Gamma_{G_{i}}^{o}(S)=\Gamma_{G_{j_{1}}}^{o}(S_{1})\oplus\Gamma_{G_{j_{2}}}^{o}(S_{2})\oplus\Gamma^{o}_{G[B_{t_{i}}]}(S_{0}).$ (35)
Taken together, (29), (30), and (35) prove the claim.
[…]
  • If $L=\emptyset$, then $S_{1}$ and $S_{2}$ may also be empty. In this situation, $f_{t_{j_{1}}}(\emptyset,\emptyset)$ and $g_{t_{j_{1}}}(\emptyset,\emptyset)$ represents the size and the number of minimum $(a,0)$ trapping sets in $G_{j_{1}}$, excluding the case of the empty set $S_{1}$. Therefore, this case requires special consideration, which leads to $f_{t_{i}}(L,I)=\min\{F_{t_{i}}(L,I),f_{t_{j_{1}}}(L,I),f_{t_{j_{2}}}(L,I)\}$ and $g_{t_{i}}(L,I)=\mathbb{1}_{f_{t_{i}}(L,I)=F_{t_{i}}(L,I)}G_{t_{i}}(L,I)+\sum_{t\in\{t_{j_{1}},t_{j_{2}}\}}\mathbb{1}_{f_{t_{i}}(L,I)=f_{t}(L,I)}g_{t}(L,I)$.
@@
For each pair $(L\subset\mathcal{L}^{B}_{i},R\subset\mathcal{R}^{B}_{i})$, it takes $O(2^{k})$ time to obtain the minimum value in (25). Therefore, the computational complexity for each join node is $O(4^{k})$, while for each other type of node, it is $O(2^{k})$. Given that a nice tree decomposition $(T,(B_{t})_{t\in T})$ of the Tanner graph $G$ contains $O(n)$ nodes in $T$, the overall time complexity of the algorithm is $O(n4^{k})$. ∎
Based on Theorem 1, if the pathwidth of a Tanner graph is known, we have the following corollary.
@@
Let $G$ be a Tanner graph with pathwidth $pw(G)=k$, corresponding to an LDPC code of length $n$. Then, the size and the number of the smallest $(a,0)$ trapping set in $G$ can be determined in time $O(n2^{k})$.
@@
The proof proceeds similarly to that of Theorem 1. However, in the case where $T$ is a path, there exists no $t\in T$ that is a join node. For all other types of nodes, the algorithm requires $O(2^{k})$ time to compute $f_{t}$ and $g_{t}$, which leads to an overall computational complexity of $O(n2^{k})$ for the algorithm. ∎
By executing the algorithm in Theorem 1 several times, we can identify an $(a,0)$ trapping set with minimum $a$ in $G$. This leads to the following corollary:
@@
Let $G$ be a Tanner graph with treewidth $tw(G)=k$, corresponding to an LDPC code of length $n$. We can identify an $(a,0)$ trapping sets with minimum $a$ in $G$ in time $O(n(a+1)4^{k})$.
@@
By repeatedly applying the algorithm from Theorem 1 as described in Algorithm 1, a minimum $(a,0)$ trapping set can be identified. The MTS algorithm is a slight modification of the algorithm presented in Theorem 1. It operates on the graph $G$ together with a tree decomposition $(T,(B_{t})_{t\in T})$, and returns a node $u$ such that, for some $t_{i}\in T$ and its child $t_{j}$, $B_{t_{i}}=B_{t_{j}}\setminus\{u\}$, and the corresponding value $f_{t_{i}}(v,\emptyset)$ reaches $a$ for the first time during the execution of the algorithm. ∎
1Let $a_{0}$ be the minimum size of an $(a,0)$ trapping set, as computed by the algorithm in Theorem 1;
2 Add a variable node $v$ to $G$ such that $v$ is not connected to any check node and appears in all $B_{t}$, $t\in T$;
3 $\mathcal{A}$ =[];
4 $a=a_{0}+1$;
5 foreach $i=1:a_{0}$ do
6 $u=MTS(G,T,(B_{t})_{t\in T},v,a)$;
7
8 $\mathcal{A}$ = $\mathcal{A},u$];
9 a=a1a=a-1;
10 if $|\mathcal{A}|=a_{0}$ then
11      break;
12
13   end if
14  Re-connect $v$, where $\Gamma_{G}^{o}(v)=\Gamma_{G}^{o}(u)\oplus\Gamma_{G}^{o}(v)$;
15   Remove the node $u$ from $G$ and all $B_{t},t\in T$;
16
17 end foreach
18Return $\mathcal{A}$;
[…]